solve for theta tan3(theta) + tan(theta) = 2tan2(theta)

Grade:11

1 Answers

Ajay
7 years ago
The Problem can be written astan3(theta) + tan(theta) - 2tan2(theta) = 0
Use following formulas in above equation so that all variables are in form of Tan (theta)
Tan 3A = tan (A + 2A) = ( Tan A + Tan 2A )/(1-Tan A Tan 2A)
Tan 2A = 2Tan A/(1 – Tan^2 A)